OpenStudy (anonymous):

Consider the function f(x) = x2 + 3x − 7. Find the value of f(x) when x = 3 and select the correct answer below. A. 5 B. 8 C. 11 D. 18

OpenStudy (anonymous):

Please help with this, I don't understand

jimthompson5910 (jim_thompson5910):

everywhere you see an 'x' replace it with 3 (since x = 3) \[\Large f(x) = x^2 + 3x - 7\] \[\Large f(\color{red}{x}) = (\color{red}{x})^2 + 3(\color{red}{x}) - 7\] \[\Large f(\color{red}{3}) = (\color{red}{3})^2 + 3(\color{red}{3}) - 7\]

jimthompson5910 (jim_thompson5910):

now either use pencil/paper or a calculator to evaluate \(\Large (3)^2+3(3)-7\)

OpenStudy (anonymous):

Thank you but I still don't know how to evaluate it afterwards, my math is a little rusty

jimthompson5910 (jim_thompson5910):

you can use pencil/paper or you can use a calculator http://web2.0calc.com/

jimthompson5910 (jim_thompson5910):

two things to keep in mind 1) type `^` to mean exponent example: 9 squared = 9^2 2) something like `3(3)` means `3 times 3`. The number next to the parenthesis implies multiplication